Prove that a sequence that diverges to +∞ (resp. −∞) is divergent


1
Expert's answer
2021-02-03T04:28:07-0500

A sequence "\{a_n\}_{n=1}^{+\infty}" that diverges to +∞ or -∞ is unbounded:

"a_n\to+\infty" means "\forall M>0\,\exist N\in\mathbb{N}\, \forall n>N a_n>M" that implies "\sup\limits_n| a_n|=+\infty"

"a_n\to-\infty" means "\forall M>0\,\exist N\in\mathbb{N}\, \forall n>N a_n<-M" that implies "\sup\limits_n| a_n|=+\infty"

A sequence "\{a_n\}_{n=1}^{+\infty}" that converges to a finite number A is bounded:

"\forall \varepsilon>0\,\exist N(\varepsilon)\in\mathbb{N}\, \forall n>N(\varepsilon) |a_n-A|<\varepsilon"

If we take "\varepsilon=1" then "\forall n>N(1) |a_n-A|<1", "|a_n|<|A|+1" and "\sup\limits_n |a_n|\leq \max\{|A|+1,|a_1|, |a_2|,\dots,|a_{N(1)}|\}<+\infty"

Therefore, any sequence that diverges to +∞ or -∞ is divergent.